A ball thrown vertically upward has a speed of 5.0 m/s at a height of 3.0 m, the mass of the ball is 1.2 kg. Find its mechanical energy.


1
Expert's answer
2021-03-17T16:55:48-0400

"E=KE+PE=\frac{mv^2}{2}+mgh=\frac{1.2\cdot 5^2}{2}+1.2\cdot 9.8\cdot 3=50.28\ (J)" . Answer